I'm doing an interest problem and I would like some help please. thanks.
Find the present value. Round to the nearest cent. To get $6,500 after 8 years at 10% compounded quarterly.

Answer by mananth(16946)

Principal P = x
Amount= 6500
years=n 8
compounded 4 times a year
Rate = 10.00 (0.100 )
Amount = P*((n+r)/n)^n

6500 = x *( 1 + r/n )^ n


6500 = P ( 1 + 0.100 / 4 ) ^ 4 * 8

6500 = P( 1.02500 )^ 32

6500 /( 1.03 )^ 32 = P

P= 2949.51